Chandra spent 2/5 of her income of January for rent, and 3/4 of the remainder on other expenses.If she put the remaining $180 in her savings account, how much was her income in January?
(A)$1,000 (B) $1,200 (C) $1,400 (D) $ 1,600 (E) None of these.

Answer by mananth(16949)

Chandra spent 2/5 of her income of January for rent, and 3/4 of the remainder on other expenses.If she put the remaining $180 in her savings account, how much was her income in January?
let the income be x
rent 2/5 *x
remaing amount = 3/4(x- 2/5*x)
2/5 *x + 3/4(x-2/5*x) +180 =x
2/5*x +3/4* (5x-2x/5) +180= x
2x/5 + 9x/20 +180=x
(8x+9x-20x)/20= -180
-3x= -3600
x= -3600/-3 = 1200
the income is $1200